A lab orders 100 rats a week. Prices for 100 rats follow the following​ distribution:

Price (x) in $ Probability (P) X^2 X*P X^2*P
15 0.39
20 0.34
25 0.27

A. The probability that the price of purchasing 100 rats is​ $20.0 is:

​B. The probability that the price of purchasing 100 rats is less than​ $21.0 is:

C. The expected price of ordering 100 rats per week is $:

D. The variance for the price of ordering 100 rats per week is:

E. The standard deviation for the price of ordering 100 rats per week is:
Answer #1

solution:

From the given information

A lab orders 100 rats a week. Prices for 100 rats follow the following​ distribution:   

Price (x) in$ Probability P(X) X^2 X*P(X) X^2 * P(X)
15 0.39 225 5.85 87.75
20 0.34 400 6.80 136
25 0.27 625 6.75 168.75
Total 19.4 392.5

A) Probability that price of purchasing 100 rats is $20 is = P(X=20) = 0.34

B)  The probability that the price of purchasing 100 rats is less than​ $21.0 = P(X<21)

= P(X=15) + P(X=20)

= 0.39 + 0.34

= 0.73

C) The expected price of ordering 100 rats per week = E{X] =\sum X*P(X) = $19.4

D) The variance for the price of ordering 100 rats per week (\sigma^2) = E[X^2]-E[X]^2

= \sum X^2*P(X) - (\sum X*P(X))^2

= 392.5 - (19.4)^2

= $16.14

E) The standard deviation for the price of ordering 100 rats per week (\sigma) = \sqrt 16.14 =~ $ 4.017